3. A motorcycle and its driver have a total mass of 485.0 kg. How fast in km/h is the motorcycle moving if it has kinetic energy of 30.0 ? 


1
Expert's answer
2021-01-10T18:28:31-0500

By the definition of the kinetic energy, we have:


"KE=\dfrac{1}{2}mv^2."

Then, from this formula, we can find the velocity of the motorcycle:


"v=\sqrt{\dfrac{2KE}{m}},""v=\sqrt{\dfrac{2\cdot 30.0\ J}{485\ kg}}=0.35\ \dfrac{m}{s}\cdot(\dfrac{1\ km}{1000\ m})\cdot(\dfrac{3600\ s}{1\ h})=1.26\ \dfrac{km}{h}."

Answer:

"v=1.26\ \dfrac{km}{h}."